OpenStudy (anonymous):

Find the product of (4x − 3y)^2. 16x^2 + 24xy + 9y^2 16x^2 − 9y^2 16x^2 − 24xy + 9y^2 16x^2 + 9y^2

OpenStudy (anonymous):

(4x − 3y) (4x − 3y) 16x^2 9y^2 how did i get 24xy

OpenStudy (anonymous):

*do

OpenStudy (anonymous):

you get -24xy by combining -12xy-12xy when you foil

OpenStudy (anonymous):

(4x-3y)(4x-3y)=16x^2+9y^2-12xy-12yx=16x^2+9y^2-24yx

OpenStudy (anonymous):

oh so 16^2+9y^2-24yx

OpenStudy (anonymous):

16x^2+9y^2-24yx

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

or you could write it like this 16x^2-24xy+9y^2 which is the way I get it when I foil but its the same thing

OpenStudy (owlcoffee):

So, to explain it, I'll make a little proof. Say: \[(a+b)^{2}\] By definition we know that: \[(a+b)^{2} = (a+b)(a+b)\] There exists a property that is called "distributive, let's apply it: \[a^{2}+ab+ba+b ^{2}\] by conmutative, we know that "ab" and "ba" are equal: So we get: \[a ^{2}+2ab+b ^{2}\] Good, but... Why did I do it? looking at the problem: \[(4x-3y)^{2}\] Hey! it has the form of what I've just proven, so we can say that a=(4x) and b=(-3y) we would end with this: \[(4x)^{2}+2(4x)(-3y)+(-3y)^{2}\] Operating it we get: \[16x ^{2}-24xy+9y ^{2}\] And you're ready to go :)

OpenStudy (anonymous):

his answer is wrong dont believe it